# Transition of Anticoncentration in Gaussian Boson Sampling

### 13 Dec 2023 v1

@@ Abstract
Broadly speaking, the hardness of Boson Sampling is based on the connection between output probabilities and the permanent, which is, classically, #$\#\mathsf{P}$-hard to compute Valiant (1979). Similarly, the hardness of GBS arises from the fact that output probabilities are controlled by a generalization of the permanent called the hafnian; while the permanent counts the number of perfect matchings in a weighted bipartite graph, the hafnian counts the number of perfect matchings in an arbitrary weighted graph. Because the hafnian generalizes the permanent, it is also difficult to compute classically.
However, the complexity of classically computing an individual output probability defined in terms of the permanent or the hafnian is not itself sufficient to prove hardness of sampling from the overall probability distributions. The standard hardness argument based on Stockmeyer’s algorithm Stockmeyer (1983); Hangleiter and Eisert (2023) requires that outcome probabilities of random Boson Sampling instances be hard to approximate. Jointly with provable hardness of nearly exactly computing output probabilities Deshpande et al. (2022a), anticoncentration of the outcome probabilities serves as evidence for this. Intuitively, if most outcome probabilities are comparable to the uniform probability, then a good classical sampling algorithm needs a very precise idea of each probability’s relative magnitude because all of them are important. Anticoncentration quantifies this idea as, most concisely, the outcome-collision probability (i.e. the probability of getting the same outcome from two independent samples) of the GBS distribution averaged over the choice of linear-optical unitary and normalized by the size of the effective sample space (Hangleiter and Eisert, 2023, Sec. D). While a weak form of anticoncentration holds in Boson Sampling Aaronson and Arkhipov (2013), under what conditions anticoncentration holds in GBS is an open question.
@@
We will work in the regime in which the output states are, with high probability, (photon-)collision-free, meaning that $n_{i}\in\{0,1\}$ for all $i$. A sufficient condition for this to be the case is that the expected number of photons $\operatorname*{\mathbb{E}}[2n]=k\sinh^{2}(r)=o(\sqrt{m})$. In this regime, Ref. Deshpande et al. (2022a) provides evidence that, for any fixed $n=o(\sqrt{m})$, the distribution over submatrices is well-captured by a generalization of the circular orthogonal ensemble (COE)  [^3] Note that, strictly speaking, the conjecture is only formulated for the regime $n\leq k$ in Ref. Aaronson and Arkhipov (2013). However, the evidence provided there for the case $k=n$—Ref. Aaronson and Arkhipov (2013) proves that $n\times n$ submatrices of Haar-random unitaries are approximately Gaussian—clearly also holds for the case $k\leq n$..
@@
We work under the assumption that Conjecture 1 is true. Fixing the measured number of photons $2n$, the normalized average (outcome-)collision probability, which quantifies anticoncentration, can be written as $|\Omega_{2n}|\operatorname*{\mathbb{E}}_{U\in U(m)}[\sum_{\mathbf{n}\in\Omega}P_{U}(\mathbf{n})^{2}]$, where $|\Omega_{2n}|$ is the size of the effective sample space. Conjecture 1 implies that, with respect to random choices of $U$, all outcomes are equally distributed over the unitaries, the so-called hiding property. This implies that the inverse size of the (effective) sample space is given by the first moment $\mathbb{E}_{U}[P_{U}(\mathbf{n})]$  [^4] What we call the effective sample space here can differ from the actual sample space because we quantify its size by the inverse first moment $1/M_{1}$ with respect to the (generalized) COE measure. But this is only approximately equal to the exact size of the sample space because the hiding property only gives us an approximate equality between the actual GBS distribution and hafnians of generalized COE matrices. Under Conjecture 1, the anticoncentration property thus reduces to computing the moments
@@
$M_{1}(k,n)=\frac{(2n)!}{(2^{n}n!)^{2}}\sum_{\tau\in S_{2n}}\sum_{\{o_{i}\}_{i=1}^{n}}^{k}\prod_{j=1}^{n}\delta_{o_{\left\lceil\frac{\tau(2j-1)}{2}\right\rceil}o_{\left\lceil\frac{\tau(2j)}{2}\right\rceil}},$ (5)
see the Supplemental Material (SM) for details.
@@
Let us briefly discuss Eq. 5. The sum over $\tau\in S_{2n}$ and the product over index $j$ come from the definition (2); the sum over the indices $o_{i}\in[k]\coloneqq\{1,2,\ldots,k\}$ is due to an expansion of $X^{\top}X$ as a matrix product. Note that, when $\tau(2j-1)$ and $\tau(2j)$ form a tuple $(2\ell-1,2\ell)$, then the Kronecker $\delta$ always equals $1$ for index $o_{\ell}$, such that summing over $o_{\ell}$ yields a factor of $k$. When $\tau(2j-1)$ and $\tau(2j)$ do not form such a tuple, we get a nontrivial relationship between the indices that decreases the number of independent degrees of freedom, thus decreasing the number of factors of $k$ in the final answer. Therefore, to evaluate this expression, one must determine the number of “free indices” over all the permutations in $S_{2n}$. We accomplish this with our graph-theoretic approach.
@@
Figure 2: Examples of graphs used to calculate the (a) first and (b) second moments of GBS outcome probabilities. (a) $G\in\mathbb{G}^{1}_{4}$. There are eight vertices $O_{1}$ to $O_{8}$ representing the index $o$ (labeled in the left column). The black (solid) edges connect only adjacent pairs, whereas the red (dashed) edges form an arbitrary perfect matching. This graph has two connected components, meaning it would contribute $k^{2}$ to the first moment. (b) $G\in\mathbb{G}_{4}^{2}$. The black (solid) edges are, from left to right, type-1, type-2, type-3, and type-4, as denoted by the gray background. This means that $z=1+0\times 4^{0}+1\times 4^{1}+2\times 4^{2}+3\times 4^{3}=229$ (this is calculated by converting 0123 from base-4 into base-10 and then adding 1 such that the final result is in $[4^{4}]$). Note that black (solid) edges stay within two adjacent columns. Red (dashed) edges form an arbitrary perfect matching in each row. This graph contributes $k^{5}$ to the second moment, as there are five connected components.
@@
Specifically, define a graph $G_{\tau}$ as follows, see Fig. 2(a). Let $G_{\tau}$ have $2n$ vertices labeled $O_{1}$ through $O_{2n}$. These upper-case vertices are not directly equivalent to the lower-case indices in Eq. 5—instead, each index $o_{j}$ splits into two vertices $O_{\ell}$ and $O_{\ell^{\prime}}$ such that $\lceil\tau(\ell)/2\rceil=j=\lceil\tau(\ell^{\prime})/2\rceil$. Let $G_{\tau}$ have a black edge between $O_{2j-1}$ and $O_{2j}$ for all $j\in[n]$, and a red edge between $O_{\ell}$ and $O_{\ell^{\prime}}$ if $\lceil\tau(\ell)/2\rceil=\lceil\tau(\ell^{\prime})/2\rceil$. These two kinds of edges mimic two types of ways that dependencies in Eq. 5 can be induced through an index $j$. Red edges identify the $\ell$ and $\ell^{\prime}$ that map to the same value via $\tau$ and the ceiling function, hence red edges identify which vertices came from the same $o$ index. Black edges identify that the images of $12j-$ 1 and $2j$ are always linked by a Kronecker $\delta$.
@@
The proof proceeds by induction over $n$, where the inductive step reduces a graph in $\mathbb{G}_{n}^{1}$ to one in $\mathbb{G}_{n-1}^{1}$ through an analysis of the red edge connected to $O_{1}$. There are two options for this red edge: either it connects to $O_{2}$, the vertex with which $O_{1}$ shares a black edge, or it attaches to some $O_{a>2}$. The former creates a connected component of size two; the latter reduces to a graph in $\mathbb{G}_{n-1}^{1}$ by merging vertices $O_{1}$, $O_{2}$, and $O_{a}$ (which does not change the number of connected components). Full details can be found in the SM.
@@
We also again use the fact that many permutations induce the same final graph. We thus define $\mathbb{G}^{2}_{n}(z)$ to be the set of graphs for the $z$th set of black edges and $\mathbb{G}^{2}_{n}\equiv\bigcup_{z\in[4^{n}]}\mathbb{G}^{2}_{n}(z)$. Because there are now three permutations associated to each graph, we obtain a degeneracy factor of $(2^{n}n!)^{3}$ and find
@@
Transition in anticoncentration.—We now use Theorems 1 and 2 in order to show that anticoncentration in GBS undergoes a transition as a function of $k$. Roughly speaking, $m_{2}(k,n)$ upper-bounds the fractional support of the outcome distribution on outcomes with probabilities larger than uniform, i.e. those most relevant to the sampling task, as we explain in detail in the SM. We speak of strong anticoncentration if $m_{2}(k,n)\geq\mathrm{const.}$. We speak of weak anticoncentration if $m_{2}(k,n)\geq 1/\mathrm{poly}(n)$. If $m_{2}(k,n)=O(1/\exp(n))$, however, we say that there is a lack of anticoncentration; in this regime, only a negligible fraction of the probabilities are nontrivial. While our definition of anticoncentration in terms of $m_{2}$ is stronger than the standard definition, it captures the essence of anticoncentration, see the SM. We show a transition between a lack of anticoncentration for $k=1$ and weak anticoncentration for $k\rightarrow\infty$ (which, of course, requires $m\to\infty$ as well).
@@
This proves the central claim of our work.
@@
Our results also connect to results on the classical simulability of GBS. The hafnian of a matrix $A$ can be computed in time exponential in the rank of $A$ Björklund et al. (2019). The absence of anticoncentration for small $k\ll n$ thus overlaps with this regime of efficient classical simulability, as the rank of $X^{\top}X$ is upper-bounded by $k$.
But does it also extend beyond this regime? While we have been able to prove the existence of this transition, our above work is not sufficient to pin down its precise location. However, we conjecture that, analogously to the hiding transition in Scattershot Boson Sampling, it occurs at $k=\omega(n^{2})$. In a companion work Ehrenberg et al. , we give evidence for this conjecture by fully analyzing the coefficients $c_{i}$ of $M_{2}(k,n)$.
@@ Acknowledgements.
In this Supplemental Material, we provide details behind many of the expressions in the main text. In particular, we derive expressions for the first and second moments of the output probabilities in the form of Kronecker $\delta$s and derive Eqs. 5, 8 and 9 of the main text. In addition, we prove Theorem 1, Theorem 2, and Lemma 1 that are presented as building blocks toward the proof of a transition in anticoncentration. We also contextualize our definition of anticoncentration with respect to the literature and use Scattershot Boson Sampling as intuition for why the transition in anticoncentration in Gaussian Boson Sampling exists.
@@
which is Eq. 8 of the main text. Removing the degeneracies induced by different permutations, and defining $\mathbb{G}^{2}_{n}(z)$ to be the set of graphs for the $z$th set of black edges and $\mathbb{G}^{2}_{n}\equiv\bigcup_{z=1}^{4^{n}}\mathbb{G}^{2}_{n}(z)$, we get a final result of
@@
We proceed by induction on $n$. Let $f(k,n)$ be the LHS of Eq. S31. For the base case $n=1$, there is only a single possible graph $G$ that has a single connected component. Thus $f(k,1)=k$. For the inductive step, which is visualized in Fig. S1, consider two subsets of $\mathbb{G}_{n}^{1}$. The first set has graphs that possess a red edge between $O_{1}$ and $O_{2}$, which means that these two vertices form their own connected component (recall that $O_{1}$ and $O_{2}$ are always connected with a black edge). Summing $k^{C(G)}$ over all graphs of this type then yields a contribution of $kf(k,n-1)$. The other subset of $\mathbb{G}_{n}^{1}$ has graphs that possess a red edge between $O_{1}$ and a vertex besides $O_{2}$, say $O_{a}$. In these graphs, the number of connected components in the graph does not change if one collapses the three vertices $O_{1}$, $O_{2}$, and $O_{a}$ into a single vertex (because they are all connected by either a black or red edge). Therefore, because there are $22n-$ 2 choices for the vertex $O_{a}$ linked to $O_{1}$ by a red edge, we get an overall contribution of $(2n-2)f(k,n-1)$ when summing $k^{C(G)}$ over these graphs.
Figure S1: Visualization of the inductive step in the proof of the first moment of the output probabilities (Theorem 1 in the main text, restated here in the SM as Theorem S1). The inductive step proceeds in two cases that are determined by the red edge that connects to the first vertex in the graph in $\mathbb{G}_{n}^{1}$. In (a), we consider the case where the first two vertices, which are linked by a black (solid) edge, are also linked by a red (dashed) edge, meaning they comprise a single connected component. This contributes a factor of $k$ times the contribution from a graph in $\mathbb{G}_{n-1}^{1}$, which comes from the remaining $22n-$ 2 vertices and their edges. (b) considers the case where the first vertex is linked via a red edge to a different vertex $O_{a}\neq O_{2}$ for which there are $22n-$ 2 choices (here $a=3$). The number of connected components does not change after identifying and combining the three vertices that are connected in this way (visualized by the blue background), meaning we again reduce down to a graph in $\mathbb{G}_{n-1}^{1}$, but this time without the multiplicative factor of $k$.
@@
Figure S2: (a) Example graph in $\mathbb{G}_{6}^{2}$ showing how to achieve an average of two connected components per set of six vertices using only type-1 and type-4 sets of edges. All vertices connected by horizontal black (solid) edges are also connected by red (dashed) edges. All type-1 vertical edges are paired off, as are type-4 vertical edges. Note that this graph would correspond to $z=1+3\times 4^{0}+0\times 4^{1}+0\times 4^{2}+0\times 4^{3}+3\times 4^{4}+3\times 4^{5}=3844$. (b) Example showing how using type-2 and type-3 black edges lead to, at most, three connected components per two sets of six vertices.

### 20 Mar 2024 v2

@@ Abstract
Broadly speaking, the hardness of Boson Sampling is based on the connection between output probabilities and the permanent, which is, classically, #$\#\mathsf{P}$-hard to compute exactly Valiant (1979). Similarly, the hardness of GBS arises from the fact that output probabilities are controlled by a generalization of the permanent called the hafnian; while the permanent counts the number of perfect matchings in a weighted bipartite graph, the hafnian counts the number of perfect matchings in an arbitrary weighted graph Barvinok (2016). Because the hafnian generalizes the permanent, it is also difficult to compute classically.
However, the complexity of classically computing an individual output probability defined in terms of the permanent or the hafnian is not itself sufficient to prove hardness of sampling from the overall probability distributions. The standard hardness argument based on Stockmeyer’s algorithm Stockmeyer (1983); Hangleiter and Eisert (2023) requires that outcome probabilities of random Boson Sampling instances be hard to approximate. Jointly with provable hardness of nearly exactly computing output probabilities Deshpande et al. (2022a), anticoncentration of the outcome probabilities serves as evidence for this. Intuitively, if most outcome probabilities are comparable to the uniform probability, then a good classical sampling algorithm needs a very precise idea of each probability’s relative magnitude because all of them are important. Anticoncentration quantifies this idea as, most concisely, the outcome-collision probability (i.e. the probability of getting the same outcome from two independent samples) of the GBS distribution averaged over the choice of linear-optical unitary and normalized by the size of the sample space (Hangleiter and Eisert, 2023, Sec. D). While a weak form of anticoncentration holds in Boson Sampling Aaronson and Arkhipov (2013), under what conditions anticoncentration holds in GBS is an open question.
@@
We will work in the regime in which the output states are, with high probability, (photon-)collision-free, meaning that $n_{i}\in\{0,1\}$ for all $i$. A sufficient condition for this to be the case is that the expected number of photons $\operatorname*{\mathbb{E}}[2n]=k\sinh^{2}(r)=o(\sqrt{m})$. In this regime, Ref. Deshpande et al. (2022a) provides evidence that, for any fixed $n=o(\sqrt{m})$, the distribution over submatrices is well-captured by a generalization of the circular orthogonal ensemble (COE)  [^3] Note that, strictly speaking, the conjecture is only formulated for the regime $n\leq k$ in Ref. Deshpande et al. (2022a). However, the evidence provided there for the case $k=n$—Ref. Aaronson and Arkhipov (2013) proves that $n\times n$ submatrices of Haar-random unitaries are approximately Gaussian—clearly also holds for the case $k\leq n$..
@@
We work under the assumption that Conjecture 1 is true. Fixing the measured number of photons $2n$, the normalized average (outcome-)collision probability, which quantifies anticoncentration, can be written as $|\Omega_{2n}|\operatorname*{\mathbb{E}}_{U\in U(m)}[\sum_{\mathbf{n}\in\Omega_{2n}}P_{U}(\mathbf{n})^{2}]$, where $|\Omega_{2n}|$ is the size of the photon-non-collisional sample space of $2n$ photons in $m$ modes, which is the dominant space of outputs when $n=o(\sqrt{m})$. Conjecture 1 implies that, with respect to random choices of $U$, all outcomes are equally distributed over the unitaries, the so-called hiding property. This implies that the inverse size of the sample space is given by the first moment $\mathbb{E}_{U}[P_{U}(\mathbf{n})]$. See Section S4 of the Supplemental Material (SM) for more details. Under Conjecture 1, the anticoncentration property thus reduces to computing the moments
@@
$M_{1}(k,n)=\frac{(2n)!}{(2^{n}n!)^{2}}\sum_{\tau\in S_{2n}}\sum_{\{o_{i}\}_{i=1}^{n}}^{k}\prod_{j=1}^{n}\delta_{o_{\left\lceil\frac{\tau(2j-1)}{2}\right\rceil}o_{\left\lceil\frac{\tau(2j)}{2}\right\rceil}}.$ (5)
@@
Let us briefly discuss Eq. 5, see the SM for details. The sum over $\tau\in S_{2n}$ and the product over index $j$ come from Eq. 2; the sum over the indices $o_{i}\in[k]\coloneqq\{1,2,\ldots,k\}$ is due to an expansion of $X^{\top}X$ as a matrix product. Note that, when $\tau(2j-1)$ and $\tau(2j)$ form a tuple $(2\ell-1,2\ell)$, then the Kronecker $\delta$ always equals $1$ for index $o_{\ell}$, such that summing over $o_{\ell}$ yields a factor of $k$. When $\tau(2j-1)$ and $\tau(2j)$ do not form such a tuple, we get a nontrivial relationship between indices that decreases the number of independent degrees of freedom, thus decreasing the number of factors of $k$ in the final answer. Therefore, to evaluate this expression, one must determine the number of “free indices” over all the permutations in $S_{2n}$. We accomplish this with our graph-theoretic approach.
@@
Figure 2: Examples of graphs used to calculate the (a) first and (b) second moments of GBS outcome probabilities. (a) $G\in\mathbb{G}^{1}_{4}$. There are eight vertices $O_{1}$ to $O_{8}$ representing the index $o$ (labeled in the left column). The black (solid) edges connect only adjacent pairs, whereas the red (dashed) edges form an arbitrary perfect matching. This graph has two connected components, meaning it contributes $k^{2}$ to the first moment. (b) $G\in\mathbb{G}_{4}^{2}$. The black (solid) edges are, from left to right, type-1, type-2, type-3, and type-4, as denoted by the gray background. $z=1+0\times 4^{3}+1\times 4^{2}+2\times 4^{1}+3\times 4^{0}=28$ (this is calculated by converting 0123 from base-4 into base-10 and then adding 1 such that the final result is in $[4^{4}]$). Note that black (solid) edges stay within two adjacent columns. Red (dashed) edges form an arbitrary perfect matching in each row. This graph contributes $k^{5}$ to the second moment, as there are five connected components.
@@
Specifically, define a graph $G_{\tau}$ as follows, see Fig. 2(a). Let $G_{\tau}$ have $2n$ vertices labeled $O_{1}$ through $O_{2n}$. These upper-case vertices are not directly equivalent to the lower-case indices in Eq. 5—instead, each index $o_{j}$ splits into two vertices $O_{\ell}$ and $O_{\ell^{\prime}}$ such that $\lceil\tau(\ell)/2\rceil=j=\lceil\tau(\ell^{\prime})/2\rceil$ (in other words, $o_{\lceil\tau(\ell)/2\rceil}$ maps to a vertex $O_{\ell}$). Let $G_{\tau}$ have a black edge between $O_{2j-1}$ and $O_{2j}$ for all $j\in[n]$, and a red edge between $O_{\ell}$ and $O_{\ell^{\prime}}$ if $\lceil\tau(\ell)/2\rceil=\lceil\tau(\ell^{\prime})/2\rceil$. These two kinds of edges mimic two types of ways that dependencies in Eq. 5 can be induced through an index $j$. Red edges identify the $\ell$ and $\ell^{\prime}$ that map to the same value via $\tau$ and the ceiling function, hence red edges identify which vertices came from the same $o$ index. Black edges identify that Eq. 5 has a Kronecker $\delta$ between $o_{\lceil{\tau(2j-1)}/2\rceil}$ and $o_{\lceil{\tau(2j)}/2\rceil}$.
@@
The proof proceeds by induction over $n$, where the inductive step reduces a graph in $\mathbb{G}_{n}^{1}$ to one in $\mathbb{G}_{n-1}^{1}$ through an analysis of the red edge connected to $O_{1}$. There are two options for this red edge: either it connects to $O_{2}$, the vertex with which $O_{1}$ shares a black edge, or it attaches to some $O_{x>2}$. The former creates a connected component of size two; the latter reduces to a graph in $\mathbb{G}_{n-1}^{1}$ by merging vertices $O_{1}$, $O_{2}$, and $O_{x}$ (which does not change the number of connected components). Full details can be found in the SM.
@@
We also again use the fact that many permutations induce the same final graph. We thus define $\mathbb{G}^{2}_{n}(z)$ to be the set of graphs for the $z$th set of black edges and $\mathbb{G}^{2}_{n}\coloneqq\bigcup_{z\in[4^{n}]}\mathbb{G}^{2}_{n}(z)$. Because there are now three permutations associated to each graph, we obtain a degeneracy factor of $(2^{n}n!)^{3}$ and find
@@
The second moment $M_{2}(k,n)$ is a degree-$2n$ polynomial in $k$ and can be written as $M_{2}(k,n)=\mbox{$(2n-1)!!$}\sum_{i=1}^{2n}c_{i}k^{i}$, where $c_{i}$ is the number of graphs $G\in\mathbb{G}^{2}_{n}$ that have $i$ connected components.
@@
Transition in anticoncentration.—We now use Theorems 1 and 2 in order to show that anticoncentration in GBS undergoes a transition as a function of $k$. Roughly speaking, $m_{2}(k,n)$ upper-bounds the fractional support of the outcome distribution on outcomes with probabilities larger than uniform, i.e. those most relevant to the sampling task, as we explain in detail in the SM. We speak of strong anticoncentration if $m_{2}(k,n)\geq\mathrm{const.}$. We speak of weak anticoncentration if $m_{2}(k,n)\geq 1/\mathrm{poly}(n)$. If $m_{2}(k,n)=O(1/n^{a})$ for any constant $a>0$, however, we say that there is a lack of anticoncentration; in this regime, only a negligible fraction of the probabilities are nontrivial. While our definition of anticoncentration in terms of $m_{2}$ is stronger than the standard definition, it captures the essence of anticoncentration, see the SM. We show a transition between a lack of anticoncentration for $k=1$ and weak anticoncentration for $k\rightarrow\infty$ (which, of course, requires $m\to\infty$ as well).
@@
This proves the central claim of our work. In the SM, we also show how anticoncentration of the approximate GBS distribution relates to anticoncentration of the true distribution.
@@
Our results also connect to the classical simulability of GBS. The hafnian of $A$ can be computed in time exponential in the rank of $A$ Björklund et al. (2019). The absence of anticoncentration for small $k\ll n$ overlaps with this regime of efficient classical simulability, as the rank of $X^{\top}X$ is upper-bounded by $k$.
But does it also extend beyond this regime? While we have been able to prove the existence of this transition, our above work is not sufficient to pin down its precise location. However, we conjecture that, weak anticoncentration holds for $k=\omega(n^{2})$, but there is a lack of anticoncentration for $k=O(n^{2})$. In a companion work Ehrenberg et al. , we give evidence for this conjecture by fully analyzing the coefficients $c_{i}$ of $M_{2}(k,n)$.
@@ Acknowledgements.
In this Supplemental Material, we provide details behind many of the expressions in the main text. In particular, we derive expressions for the first and second moments of the output probabilities in the form of Kronecker $\delta$s and derive Eqs. 5, 8 and 9 of the main text. In addition, we prove Theorem 1, Theorem 2, and Lemma 1 that are presented as building blocks toward the proof of a transition in anticoncentration. We also explain more thoroughly the connection between the hiding property and the first moment of hafnians of generalized COE matrices. We further contextualize our definition of anticoncentration with respect to the literature and show how anticoncentration of the approximate distribution of output probabilities connects to anticoncentration of the true distribution. Finally, we use Scattershot Boson Sampling as intuition for why the transition in anticoncentration in Gaussian Boson Sampling exists.
@@
which is Eq. 8 of the main text. Removing the degeneracies induced by different permutations, and defining $\mathbb{G}^{2}_{n}(z)$ to be the set of graphs for the $z$th set of black edges and $\mathbb{G}^{2}_{n}\coloneqq\bigcup_{z=1}^{4^{n}}\mathbb{G}^{2}_{n}(z)$, we get a final result of
@@
We proceed by induction on $n$. Let $f(k,n)$ be the LHS of Eq. S31. For the base case $n=1$, there is only a single possible graph $G$ that has a single connected component. Thus $f(k,1)=k$. For the inductive step, which is visualized in Fig. S1, consider two subsets of $\mathbb{G}_{n}^{1}$. The first set has graphs that possess a red edge between $O_{1}$ and $O_{2}$, which means that these two vertices form their own connected component (recall that $O_{1}$ and $O_{2}$ are always connected with a black edge). Summing $k^{C(G)}$ over all graphs of this type then yields a contribution of $kf(k,n-1)$. The other subset of $\mathbb{G}_{n}^{1}$ has graphs that possess a red edge between $O_{1}$ and a vertex besides $O_{2}$, say $O_{x}$. In these graphs, the number of connected components in the graph does not change if one collapses the three vertices $O_{1}$, $O_{2}$, and $O_{x}$ into a single vertex (because they are all connected by either a black or red edge). Therefore, because there are $22n-$ 2 choices for the vertex $O_{x}$ linked to $O_{1}$ by a red edge, we get an overall contribution of $(2n-2)f(k,n-1)$ when summing $k^{C(G)}$ over these graphs.
Figure S1: Visualization of the inductive step in the proof of the first moment of the output probabilities (Theorem 1 in the main text, restated here in the SM as Theorem S1). The inductive step proceeds in two cases that are determined by the red edge that connects to the first vertex in the graph in $\mathbb{G}_{n}^{1}$. In (a), we consider the case where the first two vertices, which are linked by a black (solid) edge, are also linked by a red (dashed) edge, meaning they comprise a single connected component. This contributes a factor of $k$ times the contribution from a graph in $\mathbb{G}_{n-1}^{1}$, which comes from the remaining $22n-$ 2 vertices and their edges. (b) considers the case where the first vertex is linked via a red edge to a different vertex $O_{x}\neq O_{2}$ for which there are $22n-$ 2 choices (here $a=3$). The number of connected components does not change after identifying and combining the three vertices that are connected in this way (visualized by the blue background), meaning we again reduce down to a graph in $\mathbb{G}_{n-1}^{1}$, but this time without the multiplicative factor of $k$.
@@
Figure S2: (a) Example graph in $\mathbb{G}_{6}^{2}$ showing how to achieve an average of two connected components per set of six vertices using only type-1 and type-4 sets of edges. All vertices connected by horizontal black (solid) edges are also connected by red (dashed) edges. All type-1 vertical edges are paired off, as are type-4 vertical edges. Note that this graph would correspond to $z=1+3\times 4^{5}+0\times 4^{4}+0\times 4^{3}+0\times 4^{2}+3\times 4^{1}+3\times 4^{0}=3088$. (b) Example showing how using type-2 and type-3 black edges lead to, at most, three connected components per two sets of six vertices.